what would be osmotic pressure of 0.05N sucrose solution at 5 degrees C ? find out the concentration of glucose solution which is isotonic with this sucrose solution.

Grade:12th pass

1 Answers

Rishita
6 years ago
P =cRTp=0.05×0.082×278atmConc. Of sucrose solution is 0.05Nor 0.05×342= 17.1g/LIsotonic solutions have same osmotic pressure so πsugar =πglucose so Csugar =C glucose 17.1/342 =unknown /180 = 9 g/L